OpenStudy (anonymous):

Find the x- and y-intercepts of the circle given by the equation . State each intercept as an ordered pair (e.g. (0,-3)), entering the smaller intercept FIRST. Give answers to the nearest tenth.

OpenStudy (anonymous):

where is the equation?

OpenStudy (anonymous):

oops, I guess it didn't come through after copy and paste. but it is (x-3)^2+(y-5)^2=100

OpenStudy (anonymous):

x intercept (x-3)^2 = 100.. square both sides u get x-3 =10.. add 3 to both sides. u get x=13

OpenStudy (anonymous):

do the same for the y value to find the y intercept

OpenStudy (anonymous):

okay makes sense, but the only problem now is that this problem asks for two values for each term

OpenStudy (anonymous):

is it asking u for 2 x intercepts, and 2 y intercepts?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that make sense.. (x-3)^2 = 100, so expand (x-3)^2 to (x-3)(x-3)=100, so the x would be x = 103 or x = 103.. do the same for y.

OpenStudy (anonymous):

okay thanks :)